# Fully non-linear elliptic equations on complex manifolds

### v2

@@ 9.1. Quantitative boundary estimate revisited
Near $z_{0}\in\partial M$, we choose the local complex coordinate (7.2) with origin at $z_{0}$ ($z=0$). In what follows, the discussion will be given at $z_{0}$. As in (7.18), we denote $\eta=(u-\underline{u})_{x^{n}}(0)$. The maximum principle tells us that $\eta\geqslant 0.$
@@ 10.2. The Dirichlet problem with less regular boundary assumption
The Dirichlet problem for homogeneous complex Monge-Ampère equation on $X\times S$ is closely connected with the geodesics in the space of Kähler potentials $\mathcal{H}_{\omega_{X}}$ [17, 47, 42]; see e.g. [4, 9, 44]. We refer to the monograph [33] for more results regarding to degenerated Monge-Ampère equation.
@@ Acknowledgements
The author was partially supported by the National Natural Science Foundation of China (Grant No. 11801587), Guangzhou Science and Technology Program (Grant No. 202201010451) and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515012121).

### v3

@@ 9.1. Quantitative boundary estimate revisited
The proof is based on Lemma 3.7. Near $z_{0}\in\partial M$, we choose the local complex coordinate (7.2) with origin at $z_{0}$ ($z=0$). In what follows, the discussion will be given at $z_{0}$. As in (7.18), we denote $\eta=(u-\underline{u})_{x^{n}}(0)$. The maximum principle tells us that $\eta\geqslant 0.$
@@ 10.2. The Dirichlet problem with less regular boundary assumption
The Dirichlet problem for homogeneous complex Monge-Ampère equation on $X\times S$ is closely connected with the geodesics in the space of Kähler potentials $\mathcal{H}_{\omega_{X}}$ [17, 47, 42]; see e.g. [4, 9, 44]. We also refer to the monograph [33] for more results regarding to degenerated Monge-Ampère equation.
@@ Acknowledgements
The author was partially supported by the National Natural Science Foundation of China (Grant No. 11801587), Guangzhou Science and Technology Program (Grant No. 202201010451), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515012121).